OpenStudy (anonymous):

What is the similarity ratio for two circles with areas 2π m2 and 200π m2?

OpenStudy (anonymous):

2π m / (200π m) = 1/100 is ratio of areas. Take square root and get 1/10 or as a ratio 1:10 that is 1 to 10 for the similarity ratio, or ratio of the radii. In other words the big circle has a radius 10 times as long as the small circle's radius.

OpenStudy (anonymous):

100:1

OpenStudy (anonymous):

so it'd be 1:100

OpenStudy (anonymous):

or Ratio of area = 2π / 200π = 2 / 200 = 1 : 100. Ratio of length = √1 / √100 = 1 : 10.

OpenStudy (anonymous):

but i can only choose one. 1:10 1:100

OpenStudy (anonymous):

hahaha....tried to proved...

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

it is 1:10

OpenStudy (anonymous):

thank you:)

OpenStudy (anonymous):

WHAT ABOUT THIS QUESTION PLEASE Cost Function is c(x) = 2x+5 and the revenue function is R(x) = 2x - 2x^2 where x is the number of units (in thousands) produced and sold and R and C are measured in millions of dollars. Find: 1. Marginal Revenue 2. Marginal Cost 3. Break-even points 4. number of x for which marginal revenue equals marginal costs.

OpenStudy (anonymous):

can you help me with another one please?

OpenStudy (anonymous):

i honestly suck at math but i recommend you ask everyone else. theyre probably a lot smarter than me, sorry. /: @PexVura

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

good luck!